OpenStudy (butterflyhope):

Urgent *will medal*: Write an equation and solve: A sugar cookie recipe calls for 3 1/2 cups of flour to make 4 dozen cookies. How much flour is needed to make 24 cookies? Thanks :)

OpenStudy (anonymous):

4 dozen is 48, so we're looking to make half quantities of this recipe. Half of 3.5 is 1.75 cups :)

OpenStudy (butterflyhope):

@tom982 how is that written as an equation?

OpenStudy (anonymous):

3.5*(24/48)=1.75

OpenStudy (ccswims):

Well 4 dozens = 48. We only need 24, which is half of that. So we only need half of the flour. f = 1/2 * 3 1/2 f = 1 3/4 of flour is needed for 24 cookies

OpenStudy (butterflyhope):

Thank you c:

OpenStudy (ccswims):

f = 1/2 (3 1/2)